Precalculus

Good day sir could you please assist for solutions to this problems for Precalculus Pre-engineering, which is due today.


*Problem 1*


For the expression

f(x)=log((1-|x|) ÷ (1+|x|))


Determine:


• it's largest domain;

• it's intersections with the x-axis (if any);

• it's intersection with the y-axis (if any);

• it's sign;

• when appropriate, end behaviours and behaviours at accumulation points of the domain which are not in the domain, possible symptoms.


*Problem 2*

Compute, showing the procedure,


limit from x to positive infinity ( square root symbol with x^2 - 3x - square root symbol with x^2 - 5x + 1)


*Problem 3*


Compute, showing the procedure.,


limit from x to negative infinity of the expression (square root with x^6 - x^2)÷(1-2x)


1
Expert's answer
2021-10-12T01:54:18-0400

Problem 1


"f(x)=\log(\dfrac{1-|x|}{1+|x|})"

"\dfrac{1-|x|}{1+|x|}>0=>1-|x|>0=>|x|<1"

"Domain: (-1, 1)"

"x-" intercept: "y=0=>0=\log(\dfrac{1-|x|}{1+|x|})=>\dfrac{1-|x|}{1+|x|}=1"



"=>1-|x|=1+|x|=>x=0"

Point "(0, 0)."

"y-" intercept: "x=0=>y(0)=\log(\dfrac{1-|0|}{1+|0|})=0"

Point "(0, 0)."

The graph passes through the origin.


"\dfrac{1-|x|}{1+|x|}=1, x=0"

"0<\dfrac{1-|x|}{1+|x|}<1, x\in(-1, 0)\cup(0, 1)"

Then "f(x)<0, x\in(-1, 0)\cup(0, 1)" and "f(0)=0."

"Range: (-\infin, \infin)"


"\lim\limits_{x\to-1^+}f(x)=\lim\limits_{x\to-1^+}\log(\dfrac{1-|x|}{1+|x|})=-\infin"

"\lim\limits_{x\to1^-}f(x)=\lim\limits_{x\to1^-}\log(\dfrac{1-|x|}{1+|x|})=-\infin"



Problem 2


"\lim\limits_{x\to\infin}(\sqrt{x^2-3x}-\sqrt{x^2-5x+1})"

"=\lim\limits_{x\to\infin}(\dfrac{x^2-3x-(x^2-5x+1)}{\sqrt{x^2-3x}+\sqrt{x^2-5x+1}})"

"=\lim\limits_{x\to\infin}(\dfrac{\dfrac{2x}{x}-\dfrac{1}{x}}{\sqrt{\dfrac{x^2}{x^2}-\dfrac{3x}{x^2}}+\sqrt{\dfrac{x^2}{x^2}-\dfrac{5x}{x^2}+\dfrac{1}{x^2}}})"

"=\lim\limits_{x\to\infin}(\dfrac{2-\dfrac{1}{x}}{\sqrt{1-\dfrac{3}{x}}+\sqrt{1-\dfrac{5}{x}+\dfrac{1}{x^2}}})"

"=\dfrac{2-0}{\sqrt{1-0}+\sqrt{1-0+0}}=1"

Problem 3


"\lim\limits_{x\to-\infin}\dfrac{\sqrt{x^6-x^2}}{1-2x}"

"=\lim\limits_{x\to-\infin}\dfrac{-x\sqrt{x^4-1}}{1-2x}"

"=\lim\limits_{x\to-\infin}\dfrac{-\sqrt{x^4-1}}{\dfrac{1}{x}-\dfrac{2x}{x}}"

"=\lim\limits_{x\to-\infin}\dfrac{\sqrt{x^4-1}}{2-\dfrac{1}{x}}=\infin"